C = 3000

I = 2000

G = 2500

T = 0.2Y

MPC = 0.5

X=6500

Z=5500 + 0.2Y

                   i.           Find the equilibrium level of income.

                 ii.           If investment expenditure decreases by 100, what will be the change in Y?

               iii.           Using the initial values, if G increases by 300 what will be the new level of Y?

               iv.           Using initial values, what will the new level of Y be if the tax rate rises to T=0.3Y?


1
Expert's answer
2021-06-20T17:49:44-0400

Part i

"y=3000+0.5(4-0.24)+2000+2500+6500-5500-0.2y"

"y=8500+0.4y-0.2y"

"0.8y=8500"

"y=10625"

Part ii

If it falls by 100

"0.8y=8500-100"

"0.8y=8400"

"y=1050"

Part iii

If it increases by 300

"0.8y=8500+300"

"0.8y=8800"

"y=1100"

Part iv

"T=0.3y"

"y=8500+0.15y"

"0.85y=8500"

"y=1000"

Therefore, G= 2500 and T = 2125

Budget deficit = G-T = 375